1. The ratio of two numbers is 3 to 5. Their sum is 136. Find the two numbers show work.
Answer by ewatrrr(24785)

 
Hi,
The ratio of two numbers is 3 to 5.
Let x represent the constant of the proportionality
3x + 5x = 136
8x = 136
x = 17
The numbers are:
3*17 = 51 and
5*17 = 85
CHECKING our answer***
51 + 85 = 136
Wish You the Best in your Studies.